OpenStudy (anonymous):

use implicit differentiation to find the derivative of 4x^3+4y^3=20xy

OpenStudy (anonymous):

12x^2+12y^2 dy/dx=20x dy/dx+20y 12x^2-20y=20xdy/dx-12y^2 dy/dx 12x^2-20y=dy/dx(20x-12y^2) dy/dx=(12x^2-20y)/(20x-12y^2)

OpenStudy (anonymous):

im confused on how to differentiate the 20xy part

myininaya (myininaya):

20 is a constant multiple just bring it down later find the derivative of (xy) using the product rule since you have a product of course

myininaya (myininaya):

the product rule is (fg)'=g*f'+f*g'

OpenStudy (anonymous):

what the derrivative of y

myininaya (myininaya):

y'